OpenStudy (anonymous):

you earn $8.50 per hour. then you receive to $9.35 per hour. find the percent increase. then find your pay per hour if you receive the same percent increase two more times.

OpenStudy (mathstudent55):

The percent increase is calculated by the following formula: \( \dfrac{ \bf new~amount - old ~amount}{\bf old ~amount} \times 100\% \)

OpenStudy (anonymous):

im learning this in school and its pretty fun and easy (: just follow the formula mathstudent55 gave you its really simple! need more help?

OpenStudy (anonymous):

so is 9.35 - 8.50 8.50 @mathstudent55

OpenStudy (anonymous):

first whats 9.35 - 8.50

OpenStudy (anonymous):

is 0.85

OpenStudy (anonymous):

ok now divide 0.85 by 8.50

OpenStudy (mathstudent55):

\(\dfrac{9.35 - 8.50}{8.50} \times 100\% = \)

OpenStudy (anonymous):

what did you get?

OpenStudy (anonymous):

0.1

OpenStudy (mathstudent55):

\( \dfrac{0.85}{8.50} \times 100 \% \) \( = 0.1 \times 100\%\) \( = 10\% \) The percent increase is 10%.

OpenStudy (anonymous):

0.1 times 100% =10% and its increase (:

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

yw

OpenStudy (mathstudent55):

Now for the second part. If you receive the same pay raise two more times, then you will end up with the following. Since now $9.35 is your salary, and the full salary is 100%, a ten pecent increase would make it 110% of what you are receiving now. 110% * $9.35 =1.1 * $9.35 = $10.285 After the first of the two 10% percent increases, the new salary is $10.285. $10.285 is now 100% of your salary. A second 10% increase means the new salary becomes 110% of the current salary, so the new salary after the second 10% increase would be: 110% * $10.285 = 1.1 * $10.285 =$11.3135 The new salary is $11.31 per hour.

OpenStudy (anonymous):

thank you so much

OpenStudy (mathstudent55):

You are welcome.